40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Dec 4 13:56:54 2011, Den, Update, IOO, WFS 
    Reply  Sun Dec 4 23:49:38 2011, Den, Update, IOO, WFS 
Message ID: 6067     Entry time: Sun Dec 4 23:49:38 2011     In reply to: 6066
Author: Den 
Type: Update 
Category: IOO 
Subject: WFS 


Yesterday I locked the MC and left at 8 pm. Analyzing the data I saw that MC was locked all time from 8 pm to 12.30 am when it lost lock. Moreover there was no light on transmition and reflected screens at all. I went to the PSL and saw that no red light comes to the MC from PSL, only green. I took infrared sensos to track the laser light. Then I came back to control room to study the medm diagram of the PSL. Then I came back and saw that the laser beam goes to the MC! I returned to control room and saw light on the MC screens. Does someone do something parallel with me through ssh?

I enabled the auto locker and saw the MC locked for a couple of seconds. After that the WFS were turned on automatically and I saw that the signal of the OSEM local sensors of the MC mirrors began to increase. So the WFS master provides not good feedback signal. I thought that it is due to my recompilation of c1mcs with a fixed if-statement line. And may be if c1mcs workes without digital noise and c1ioo with it then there might occur some mismatches and the signal is corrupted. For this assumption I've recompiled c1mcs back to 1e-20 in the if-statement and so added the digital noise back that I saw in the dtt tools.

However, the problem was still present - WFS feedback signal crashed the MC lock. I open the WFS master window and disabled the output to MC. I can see that the C1:IOO-WFS1_PIT_INMON and other input channels have reasonable values 8 - 20 but the output continues to increase up to 1000000. The output was off so the MC stayed at lock. As for now I turned off WFS so no feedback is applied to MC mirros.

With the help of Suresh we have adjusted optics near PMC and input to the MC on the PSL and in the black box where WFS are. Surprisingly, some optics near WFS was not attached to the table. But these mirrors are not used. One screw was near the hole but not screwed in. This mirror is used. Suresh could also rotate other screws. I thought that they must be attached to the table more rigidly.

Then we found that WFS output matrix is wrong and Suresh recalculated it. After that we've locked the MC using WFS. C1:IOO-MC_RFPD_DCMON is 0.7-0.8. 

We also recompiled and reinstalled C1MCS and C1IOO with fixed if-statement and again saw how MC_F curve moves down. WFS error signals are also improved. But still some more work on output matrix is needed.

ELOG V3.1.3-